40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jun 25 06:09:38 2011, kiwamu, Update, LSC, Friday locking 
    Reply  Tue Jun 28 15:20:08 2011, kiwamu, Update, LSC, sensing matrix measurement in PRMI configuration rawmatrix.pngsensmat.png
       Reply  Wed Jun 29 12:20:53 2011, kiwamu, Update, LSC, sensing matrix measurement in PRMI configuration 
Message ID: 4884     Entry time: Sat Jun 25 06:09:38 2011     Reply to this: 4899
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Friday locking 

I was able to measure the sensing matrix in the PRMI configuration.

The results will be posted later.

ELOG V3.1.3-